OpenStudy (anonymous):

\[\sqrt[3]{-128}\] Help me solve!

OpenStudy (anonymous):

I would divide 128 by 2 and keep dividingit by 2 until u can no longer divideit by 2. Then each triad of 2's is one factor of the cube root. Collect all your triad of 2's.

OpenStudy (anonymous):

\[128\] is a power of 2, it is \[128=2^7\]

OpenStudy (anonymous):

I multiplied 5 three times and ended up with 125

OpenStudy (anonymous):

How would I find \[\sqrt[3]{-128}\]

OpenStudy (e.mccormick):

First, when taking a nagative out of a root, that can happen in cubes. Now, with a cube root, you have to take them out in sets of three. As satellite73 pointed out, this is \(2^7\) so you really have: \(\sqrt[3]{-(2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 2)} \). How many sets of three, 2s can you get out of that?

OpenStudy (anonymous):

2 sets.

OpenStudy (e.mccormick):

OK. What would the two sets become outside the radical?

OpenStudy (anonymous):

4?

OpenStudy (e.mccormick):

Yes! So now you have something with a 4 and a cube root of 2. The question remainging is the negative sign.

OpenStudy (e.mccormick):

Becuase it is a cube root, what do you think happens with the negative?

OpenStudy (anonymous):

It becomes a -4?

OpenStudy (e.mccormick):

Yes. However, the - inside the cube goes away. \(-4\sqrt[3]{2}\) is the best answer. \(4\sqrt[3]{-2}\) is an OK answer because it still keeps the negative, but not how it would usually be seen. \(-4\sqrt[3]{-2}\) is a BAD answer and wrong because it has two negatives and therefore is positive.

OpenStudy (e.mccormick):

And if you put \(-4\sqrt[3]{2}\) and \(\sqrt[3]{-128}\) in a calculator, you should get the same decimal aproximation answer for both!